OpenStudy (annie12m):

A pan contained 3/4 of a pizza. Jack ate 1/2 of the remaining pizza. How much of the whole pizza did Jack eat?

OpenStudy (annie12m):

3/8??

OpenStudy (annie12m):

@Ray1998 @RhondaSommer @safyousuff97

OpenStudy (anitas.):

I don't know if this is to late or not but this is how you do it: 1. The amount of pizza left in the pan = 3/4 Jack ate 1/2 of the pizza that was left in the pan. 2. Amount of pizza ate by Jack = (3/4) * (1/2) = 3/8 Now it is required to find the amount of pizza ate by Jack from the whole pizza. Then 3. Amount of pizza that Jack ate from the whole = 1 - (3/8) = (8-3)/8 = 5/8 4. So Jack ate 5/8 part of the total pizza. I hope this is the answer you were looking for. I have simplified the procedure for your understanding. Hope it has come to your help.

OpenStudy (anitas.):

@annie12m